# Learning Temporal 3D Semantic Scene Completion via Optical Flow Guidance

### v1

@@ Abstract
3D Semantic Scene Completion (SSC) provides comprehensive scene geometry and semantics for autonomous driving perception, which is crucial for enabling accurate and reliable decision-making. However, existing SSC methods are limited to capturing sparse information from the current frame or naively stacking multi-frame temporal features, thereby failing to acquire effective scene context. These approaches ignore critical motion dynamics and struggle to achieve temporal consistency. To address the above challenges, we propose a novel temporal SSC method FlowScene: Learning Temporal 3D Semantic Scene Completion via Optical Flow Guidance. By leveraging optical flow, FlowScene can integrate motion, different viewpoints, occlusions, and other contextual cues, thereby significantly improving the accuracy of 3D scene completion. Specifically, our framework introduces two key components: (1) a Flow-Guided Temporal Aggregation module that aligns and aggregates temporal features using optical flow, capturing motion-aware context and deformable structures; and (2) an Occlusion-Guided Voxel Refinement module that injects occlusion masks and temporally aggregated features into 3D voxel space, adaptively refining voxel representations for explicit geometric modeling. Experimental results demonstrate that FlowScene achieves state-of-the-art performance on the SemanticKITTI and SSCBench-KITTI-360 benchmarks.
@@ 1 Introduction
One of the key challenges in autonomous driving is 3D scene understanding, which involves interpreting the spatial layout and semantic properties of objects within the scene. The ability to perceive and accurately interpret 3D scenes is essential for making safe and informed driving decisions. Recently, the 3D Semantic Scene Completion (SSC) task [31, 30] has gained significant attention in autonomous driving, as it enables the joint inference of geometry and semantics from incomplete observations.
@@
Most existing SSC methods [29, 49, 7, 17, 30, 44] rely on input RGB images along with corresponding 3D data to predict volume occupancy and assign semantic labels. However, the dependence on 3D data often requires specialized and costly depth sensors, which can limit the broader applicability of SSC algorithms. Recently, many researchers [3, 50, 14, 10] have investigated camera-based approaches to reconstruct dense 3D geometric structures and recover semantic information, offering a more accessible alternative.
@@
Previous camera-based SSC methods [13, 47, 14] typically rely on the limited observations available in the current frame to recover 3D geometry and semantics. Later, some researchers [20, 15, 25, 39] stacked historical temporal features or aligned features with estimated camera poses to enrich contextual information, as shown in Figure 2(a). However, these direct temporal modeling methods overlook the scene motion context, fail to achieve temporal consistency, and inherently limit the increase of effective contextual cues. Based on these limitations, we asked: How can we accurately identify the correlation between historical frames and the current frame to guide temporal SSC modeling?
@@
In this paper, we propose a novel temporal SSC method: FlowScene, Learning Temporal 3D Semantic Scene Completion via Optical Flow Guidance. As shown in Figure 2(b), FlowScene uses optical flow to guide temporal modeling, injecting various types of information into the SSC model, such as motion, different viewpoints, deformation, texture, geometric structure, lighting, and occlusion. As shown in Figure 1, the corresponding optical flow and occlusion masks are generated from the historical and current frame images, allowing for the further derivation of scene geometry and semantic structure. The positions and semantics of the car, tree trunk, vegetation, and pole within the red box in Figure 1 are more accurate, even when they are mutually occluded. Specifically, we introduce the Flow-Guided Temporal Aggregation module to effectively enhance temporal and motion cues by incorporating motion and contextual information from previous frames. Furthermore, we design the Occlusion-Guided Voxel Refinement module, which leverages aggregated features and occlusion masks to refine 3D voxel predictions for explicit geometric modeling. To evaluate the performance of FlowScene, we conduct thorough experiments on SemanticKITTI [1] and SSCBench-KITTI360 [21, 19]. Our method achieves state-of-the-art performance. The main contributions of our work are summarized as follows:
@@ 2 Related Work
Semantic Scene Completion (SSC) aims to jointly predict the geometry and semantics of a scene in 3D space, addressing the challenges of both scene completion and semantic segmentation. SSCNet [31] was the first to introduce this task, inferring both occupancy and semantic labels from incomplete visual observations. Subsequent studies have leveraged explicit 3D representations, such as depth maps, occupancy grids, and point clouds [29, 49, 42], which provide rich geometric cues. Meanwhile, multi-modal approaches [2, 16] have integrated RGB imagery with depth information to enhance texture representation. Additionally, several works [44, 45, 24] have explored the interdependence between semantic segmentation and scene completion.
With the advent of cost-effective, vision-based autonomous driving solutions, monocular SSC methods have gained traction. MonoScene [3] was the first to infer dense 3D semantics from a single RGB image. TPVFormer [10] introduced a tri-perspective view (TPV) representation, extending BEV with two vertical planes. OccFormer [50] proposed a dual-path transformer to encode voxel features, while VoxFormer [20] introduced a two-stage pipeline for voxelized semantic scene understanding. SurroundOcc [41] employed 3D convolutions for progressive voxel upsampling and dense SSC ground truth generation. OctOcc [27] utilized an octree-based representation for semantic occupancy prediction, while NDCScene [46] redefined spatial encoding by mapping 2D feature maps to normalized device coordinates (NDC) rather than world space. MonoOcc [51] enhanced 3D volumetric representations using an image-conditioned cross-attention mechanism. H2GFormer [40] introduced a progressive feature reconstruction strategy to propagate 2D information across multiple viewpoints. Symphonize [13] extracted high-level instance features to serve as key-value pairs for cross-attention. HASSC [39] proposed a self-distillation framework to improve the performance of VoxFormer. Stereo-based methods, such as BRGScene [14], leveraged stereo depth estimation to resolve geometric ambiguities. MixSSC [38] fused forward projection sparsity with the denseness of depth-prior backward projection. CGFormer [47] utilized a context-aware query generator to initialize context-dependent queries tailored to individual input images, effectively capturing their unique characteristics and aggregating information within the region of interest. HTCL [15] decomposed temporal context learning into two hierarchical steps: cross-frame affinity measurement and affinity-based dynamic refinement.
@@
Optical flow estimation, a fundamental task in computer vision, aims to establish dense pixel-wise correspondences between consecutive frames. FlowNet [4, 12] introduced the first CNN-based end-to-end flow estimation pipeline, leveraging a hierarchical pyramid structure. PWC-Net [33] further refined this approach by incorporating multi-stage warping to handle large-displacement motion. RAFT [36] introduced an iterative, recurrent architecture that refines residual flow predictions in a fully convolutional manner. GMFlow [43] reframed optical flow as a global matching problem, directly computing feature similarities to establish correspondences.
Beyond motion estimation, optical flow has been leveraged to enhance various vision tasks. FlowTrack [53] used optical flow to enrich feature representations and improve tracking accuracy. FGFA [52] employed flow-guided feature aggregation for end-to-end video object detection. LoSh [48] utilized flow-based warping to propagate annotations across temporal neighbors, thereby boosting referring video object segmentation. DATMO [32] introduced a moving object detection and tracking framework tailored for autonomous vehicles. DeVOS [5] incorporated optical flow into scene motion modeling, using it as a prior for learnable offsets in video segmentation.
@@ 3.1 Preliminary
Given a set of RGB images $I=\{I_{t-i}\}_{i=0}^{n}$, where $n$ is the number of historical temporal images, the objective is to jointly infer the geometry and semantics of a 3D scene. This scene is represented as a voxel grid $Y\in\mathbb{R}^{X\times Y\times Z\times(M+1)}$, where $X,Y,Z$ represent the height, width, and depth in 3D space, respectively. Each voxel in the grid is assigned a unique semantic label from the set $C\in\{C_{0},C_{1},...,C_{M}\}$, where $C_{0}$ represents empty space and the remaining classes $\{C_{1},...,C_{M}\}$ correspond to specific semantic categories. Here, $M$ denotes the total number of semantic classes. The goal is to learn a transformation $Y=\theta(I_{s})$ that closely approximates the ground truth $\hat{Y}$.
@@ 3.2 Overview
We illustrate our method in Figure 3. First, we use the image encoder RepViT [37] and FPN [22] to extract the current image features, $F_{t}$, and the historical temporal features, $F_{temp}=\{F_{t-i}\}_{i=1}^{n}$. We then apply the pre-trained optical flow estimation model [43] (Section 3.3) to generate bidirectional optical flow, $Flow=\{Flow_{fwd}^{t-i\rightarrow t},Flow_{bwd}^{t-i\rightarrow t}\}_{i=1}^{n}$. The historical temporal features, $F_{temp}$, are warped using $Flow_{bwd}$ to obtain $F_{warp}=\{F_{warp}^{t-i\rightarrow t}\}_{i=1}^{n}$. The bidirectional optical flow is then used for occlusion detection through a forward and backward consistency check to obtain the cumulative mask, $M\in{0,1}^{1\times h\times w}$. Subsequently, $F_{warp}$, $F_{t}$, and $M$ are passed into the FGTA module (Section 3.4) to perform optical flow-guided temporal feature aggregation in the 2D image feature space, resulting in the aggregated feature $F_{agg}$. Next, we apply the LSS view transformation [28] to project $F_{t}$, $F_{agg}$, and $M$ into the 3D voxel space, obtaining $V_{t}$, $V_{agg}$, and $V_{mask}$, respectively. In the subsequent OGVR module (Section 3.5), the two voxel features are fused based on the occlusion information, yielding the refined voxel features, $V_{fine}$. Finally, $V_{fine}$ passes through the voxel encoder, then undergoes upsampling and linear projection to output the dense semantic voxels, $Y$.
@@ 3.3 Optical Flow Estimation
Given a reference image frame $I_{t}$ and historical frames $I_{t-i}$, the flow field $Flow^{t\rightarrow t-i}=\mathcal{F}(I_{t},I_{t-i})$ is estimated by a flow network $\mathcal{F}$ (e.g., GMFlow [43]). The feature map from the historical frame is warped to the reference frame according to the flow. The warping function is defined as
@@
where $\mathcal{W}arp(\cdot)$ is a bilinear warping function applied to all locations of each channel in the feature map, and $F_{warp}^{t-i\rightarrow t}$ represents the feature map warped from frame $t-i$ to $t$.
@@
First, we note that there is relative motion between almost all frames in an autonomous driving scenario, which results in pixels in the current image that do not have corresponding matching pixels in the historical frames; these are referred to as occluded areas. To detect occlusion, as shown in Figure 4, we use the commonly employed forward and backward consistency check technique [34, 26], which is implemented as:
$M=\mathcal{CC}(Flow_{bwd},Flow_{fwd}),$ (2)
where $\mathcal{CC}(\cdot)$ denotes the forward and backward consistency check function. For non-occluded pixels, the forward optical flow should be the inverse of the backward optical flow of the corresponding pixel in the second frame. A pixel is marked as occluded if the mismatch between the two flows exceeds a predefined threshold. Thus, we define the occlusion flag as 1 whenever the constraint is violated and 0 otherwise.
@@ 3.4 Flow-Guided Temporal Aggregation
Specifically, guided by optical flow, the historical frame features are warped to the reference frame. Features from different frames provide multiple information for the same object instance, such as motion, different viewpoints, deformations, textures, geometric structures, various lighting and occlusions. First, we assign different weights to different spatial locations, while ensuring that the spatial weights remain the same across all feature channels. At position P, if the warped feature $F_{warp}^{t-i\rightarrow t}(\text{P})$ is close to the feature $F_{t}(\text{P})$, it is assigned a larger weight. Otherwise, a smaller weight is assigned. Inspired by FGFA [52], we use the cosine similarity [23] to measure the similarity between the warped features and the reference frame features:
@@
The non-occluded regions in the historical frames usually have richer texture and feature information, which may be missing in the current frame due to visual occlusion. To address this, we enhance the current frame features, we effectively fuse spatiotemporal information through the neighborhood cross-attention mechanism  [8]. First, we select reliable non-occluded region features in the historical frames based on the occlusion mask. The reference features $F_{t}$ are used as query, and the warp features $F_{warp}$ of the non-occluded regions serve as key and value. The specific operations are as follows:
$F_{t}=\operatorname{NCA}(F_{t},(1-M)\cdot F_{warp}),$ (5)
where $\operatorname{NCA}(\cdot)$ is the neighborhood cross attention mechanism. After these operations, $F_{t}$ fuses the non-occluded region information from both the current and historical frames, providing more stable and accurate features that enhance the perception of dynamic scenes and occluded regions.

### v2

@@ Abstract
3D Semantic Scene Completion (SSC) provides comprehensive scene geometry and semantics for autonomous driving perception, which is crucial for enabling accurate and reliable decision-making. However, existing SSC methods are limited to capturing sparse information from the current frame or naively stacking multi-frame temporal features, thereby failing to acquire effective scene context. These approaches ignore critical motion dynamics and struggle to achieve temporal consistency. To address the above challenges, we propose a novel temporal SSC method FlowScene: Learning Temporal 3D Semantic Scene Completion via Optical Flow Guidance. By leveraging optical flow, FlowScene can integrate motion, different viewpoints, occlusions, and other contextual cues, thereby significantly improving the accuracy of 3D scene completion. Specifically, our framework introduces two key components: (1) a Flow-Guided Temporal Aggregation module that aligns and aggregates temporal features using optical flow, capturing motion-aware context and deformable structures; and (2) an Occlusion-Guided Voxel Refinement module that injects occlusion masks and temporally aggregated features into 3D voxel space, adaptively refining voxel representations for explicit geometric modeling. Experimental results demonstrate that FlowScene achieves state-of-the-art performance, with mIoU of 17.70 and 20.81 on the SemanticKITTI and SSCBench-KITTI-360 benchmarks.
@@ 1 Introduction
One of the key challenges in autonomous driving is 3D scene understanding, which involves interpreting the spatial layout and semantic properties of objects within the scene. The ability to perceive and accurately interpret 3D scenes is essential for making safe and informed driving decisions. Recently, the 3D Semantic Scene Completion (SSC) task song2017semantic; roldao2020lmscnet has gained significant attention in autonomous driving, as it enables the joint inference of geometry and semantics from incomplete observations.
@@
Most existing SSC methods rist2021semantic; zhang2018efficient; guo2018view; li2020aicnet; roldao2020lmscnet; yan2021sparse rely on input RGB images along with corresponding 3D data to predict volume occupancy and assign semantic labels. However, the dependence on 3D data often requires specialized and costly depth sensors, which can limit the broader applicability of SSC algorithms. Recently, many researchers cao2022monoscene; zhang2023occformer; li2023stereoscene; huang2023tri have investigated camera-based approaches to reconstruct dense 3D geometric structures and recover semantic information, offering a more accessible alternative.
@@
Previous camera-based SSC methods jiang2024symphonize; CGFormer; li2023stereoscene typically rely on the limited observations available in the current frame to recover 3D geometry and semantics. Later, some researchers li2023voxformer; li2024htcl; mei2024sgn; wang2024HASSC stacked historical temporal features or aligned features with estimated camera poses to enrich contextual information, as shown in Figure 2(a). However, these direct temporal modeling methods overlook the scene motion context, fail to achieve temporal consistency, and inherently limit the increase of effective contextual cues. Based on these limitations, we asked: How can we accurately identify the correlation between historical frames and the current frame to guide temporal SSC modeling?
@@
In this paper, we propose a novel temporal SSC method: FlowScene, Learning Temporal 3D Semantic Scene Completion via Optical Flow Guidance. As shown in Figure 2(b), FlowScene uses optical flow to guide temporal modeling, injecting various types of information into the SSC model, such as motion, different viewpoints, deformation, texture, geometric structure, lighting, and occlusion. As shown in Figure 1, the corresponding optical flow and occlusion masks are generated from the historical and current frame images, allowing for the further derivation of scene geometry and semantic structure. The positions and semantics of the car, tree trunk, vegetation, and pole within the red box in Figure 1 are more accurate, even when they are mutually occluded. Specifically, we introduce the Flow-Guided Temporal Aggregation module to effectively enhance temporal and motion cues by incorporating motion and contextual information from previous frames. Furthermore, we design the Occlusion-Guided Voxel Refinement module, which leverages aggregated features and occlusion masks to refine 3D voxel predictions for explicit geometric modeling. To evaluate the performance of FlowScene, we conduct thorough experiments on SemanticKITTI behley2019semantickitti and SSCBench-KITTI360 Liao2022kitti360; li2023sscbench. Our method achieves state-of-the-art performance. The main contributions of our work are summarized as follows:
@@ 2 Related Work
The vision-based 3D Semantic Scene Completion (SSC) solution has received widespread attention in the field of autonomous driving perception. MonoScene cao2022monoscene was the first to infer dense 3D semantics from a single RGB image. TPVFormer huang2023tri introduced a tri-perspective view (TPV) representation, extending BEV with two vertical planes. OccFormer zhang2023occformer proposed a dual-path transformer to encode voxel features, while VoxFormer li2023voxformer introduced a two-stage pipeline for voxelized semantic scene understanding. SurroundOcc wei2023surroundocc employed 3D convolutions for progressive voxel upsampling and dense SSC ground truth generation. OctOcc ouyang2024octocc utilized an octree-based representation for semantic occupancy prediction, while NDCScene yao2023ndc redefined spatial encoding by mapping 2D feature maps to normalized device coordinates (NDC) rather than world space. MonoOcc zheng2024monoocc enhanced 3D volumetric representations using an image-conditioned cross-attention mechanism. H2GFormer wang2024h2gformer introduced a progressive feature reconstruction strategy to propagate 2D information across multiple viewpoints. Symphonize jiang2024symphonize extracted high-level instance features to serve as key-value pairs for cross-attention. HASSC wang2024HASSC proposed a self-distillation framework to improve the performance of VoxFormer. Stereo-based methods, such as BRGScene li2023stereoscene, leveraged stereo depth estimation to resolve geometric ambiguities. MixSSC wang2025mixssc fused forward projection sparsity with the denseness of depth-prior backward projection. CGFormer CGFormer utilized a context-aware query generator to initialize context-dependent queries tailored to individual input images, effectively capturing their unique characteristics and aggregating information within the region of interest. HTCL li2024htcl decomposed temporal context learning into two hierarchical steps: cross-frame affinity measurement and affinity-based dynamic refinement. VLScene wang2025vlscene leveraged vision-language models to extract high-level semantic priors for SSC.
@@
Optical flow estimation, a fundamental task in computer vision, aims to establish dense pixel-wise correspondences between consecutive frames. FlowNet dosovitskiy2015flownet; ilg2017flownet2 introduced the first CNN-based end-to-end flow estimation pipeline, leveraging a hierarchical pyramid structure. PWC-Net sun2018pwc further refined this approach by incorporating multi-stage warping to handle large-displacement motion. RAFT teed2020raft introduced an iterative, recurrent architecture that refines residual flow predictions in a fully convolutional manner. GMFlow xu2022gmflow reframed optical flow as a global matching problem, directly computing feature similarities to establish correspondences. Beyond motion estimation, optical flow has been leveraged to enhance various vision tasks. FlowTrack zhu2018flowtrack used optical flow to enrich feature representations and improve tracking accuracy. FGFA zhu2017fgfa employed flow-guided feature aggregation for end-to-end video object detection. LoSh Yuan_2024_LoSh utilized flow-based warping to propagate annotations across temporal neighbors, thereby boosting referring video object segmentation. DATMO sormoli2024optical introduced a moving object detection and tracking framework tailored for autonomous vehicles. DeVOS Fedynyak_2024_DeVos incorporated optical flow into scene motion modeling, using it as a prior for learnable offsets in video segmentation.
@@ 3.1 Preliminary
Given a set of RGB images $I=\{I_{t-i}\}_{i=0}^{n}$, where $n$ is the number of historical temporal images, the objective is to jointly infer the geometry and semantics of a 3D scene. This scene is represented as a voxel grid $\mathbb{Y}\in\mathbb{R}^{X\times Y\times Z\times(M+1)}$, where $X,Y,Z$ represent the height, width, and depth in 3D space, respectively. Each voxel in the grid is assigned a unique semantic label from the set $C\in\{C_{0},C_{1},...,C_{M}\}$, where $C_{0}$ represents empty space and the remaining classes $\{C_{1},...,C_{M}\}$ correspond to specific semantic categories. Here, $M$ denotes the total number of semantic classes. The goal is to learn a transformation $\mathbb{Y}=\theta(I_{s})$ that closely approximates the ground truth $\hat{\mathbb{Y}}$.
@@ 3.2 Overview
We illustrate our method in Figure 3. First, we use the lightweight image encoder RepViT wang2023repvit and FPN lin2017feature to extract the current image features, $F_{t}$, and the historical temporal features, $F_{temp}=\{F_{t-i}\}_{i=1}^{n}$. We then apply the pre-trained optical flow estimation model xu2022gmflow (Section 3.3) to generate bidirectional optical flow, $Flow=\{Flow_{fwd}^{t-i\rightarrow t},Flow_{bwd}^{t-i\rightarrow t}\}_{i=1}^{n}$. The historical temporal features, $F_{temp}$, are warped using $Flow_{bwd}$ to obtain $F_{warp}=\{F_{warp}^{t-i\rightarrow t}\}_{i=1}^{n}$. The bidirectional optical flow is then used for occlusion detection through a forward and backward consistency check to obtain the cumulative mask, $M\in{0,1}^{1\times h\times w}$. Subsequently, $F_{warp}$, $F_{t}$, and $M$ are passed into the FGTA module (Section 3.4) to perform optical flow-guided temporal feature aggregation in the 2D image feature space, resulting in the aggregated feature $F_{agg}$. Next, we apply the LSS view transformation philion2020lift to project $F_{t}$, $F_{agg}$, and $M$ into the 3D voxel space, obtaining $V_{t}$, $V_{agg}$, and $V_{mask}$, respectively. In the subsequent OGVR module (Section 3.5), the two voxel features are fused based on the occlusion information, yielding the refined voxel features, $V_{fine}$. Finally, $V_{fine}$ passes through the voxel encoder, then undergoes upsampling and linear projection to output the dense semantic voxels, $\mathbb{Y}$.
@@ 3.3 Optical Flow Estimation
Given a reference image frame $I_{t}$ and historical frames $I_{t-i}$, the flow field $Flow^{t\rightarrow t-i}=\mathcal{F}(I_{t},I_{t-i})$ is estimated by a flow network $\mathcal{F}$ (e.g., GMFlow xu2022gmflow). The feature map from the historical frame is warped to the reference frame according to the flow. The warping function is defined as
@@
where $\mathcal{W}arp(\cdot)$ is a bilinear warping function applied to all locations of each channel in the feature map, and $F_{warp}^{t-i\rightarrow t}$ represents the feature map warped from frame tit-i to $t$.
@@
First, we note that there is relative motion between almost all frames in an autonomous driving scenario, which results in pixels in the current image that do not have corresponding matching pixels in the historical frames; these are referred to as occluded areas. To detect occlusion, as shown in Figure 5, we use the commonly employed forward and backward consistency check technique sundaram2010dense; meister2018unflow, which is implemented as:
$M=\mathcal{CC}(Flow^{t\rightarrow t-i},Flow^{t-i\rightarrow t}),$ (2)
where $\mathcal{CC}(\cdot)$ denotes the forward and backward consistency check function, and we have included a detailed explanation in the Technical Appendix. For non-occluded pixels, the forward optical flow should be the inverse of the backward optical flow of the corresponding pixel in the second frame. A pixel is marked as occluded if the mismatch between the two flows exceeds a predefined threshold. Thus, we define the occlusion flag as 1 whenever the constraint is violated and 0 otherwise.
@@ 3.4 Flow-Guided Temporal Aggregation
Specifically, guided by optical flow, the historical frame features are warped to the reference frame. Features from different frames provide multiple information for the same object instance, such as motion, different viewpoints, deformations, textures, geometric structures, various lighting and occlusions. First, we assign different weights to different spatial locations, while ensuring that the spatial weights remain the same across all feature channels. At position P, if the warped feature $F_{warp}^{t-i\rightarrow t}(\text{P})$ is close to the feature $F_{t}(\text{P})$, it is assigned a larger weight. Otherwise, a smaller weight is assigned. Inspired by FGFA zhu2017fgfa, we use the cosine similarity luo2018cosine to measure the similarity between the warped features and the reference frame features:
@@
The non-occluded regions in the historical frames usually have richer texture and feature information, which may be missing in the current frame due to visual occlusion. To address this, we enhance the current frame features, we effectively fuse spatiotemporal information through the neighborhood cross-attention mechanism  hassani_2023_neighborhood. We selectively use non-occluded features from historical frames to prevent injecting unreliable or distorted information caused by occlusions or inaccurate flow. Including occluded regions can introduce noise and harm completion quality. First, we select reliable non-occluded region features in the historical frames based on the occlusion mask. The reference features $F_{t}$ are used as query, and the warp features $F_{warp}$ of the non-occluded regions serve as key and value. The specific operations are as follows:
$F_{t}=\mathcal{NCA}(F_{t},(1-M)\cdot F_{warp}),$ (5)
where $\mathcal{NCA}(\cdot)$ is the neighborhood cross attention mechanism. After these operations, $F_{t}$ fuses the non-occluded region information from both the current and historical frames, providing more stable and accurate features that enhance the perception of dynamic scenes and occluded regions.